OpenStudy (anonymous):

Find the angle that maximizes the area of the isosceles triangle whose legs have length l=41

Directrix (directrix):

I'm looking at the formula for area of any triangle as A = one-half the product of any two sides and the sine of the included angle. A = (1/2) (41)(41) sin (theta) To maximize area, what value of theta gives sin(theta) its maximum value. That's the answer.

Directrix (directrix):

The sine function varies between -1 and 1 so what angle makes sine = 1.

OpenStudy (cwrw238):

area = 0.5 * 41^2 * sin x differentiate and equate to zero solve to get maximum area angle x

OpenStudy (anonymous):

If I diff. the area formula A=SIn(x)*(1/2)41^2 and set equal to zero x=pi

OpenStudy (anonymous):

I'm on WebWork hw site. Tried that as the answer but returned it as incorrect

OpenStudy (anonymous):

x = ± pi/2

OpenStudy (cwrw238):

yup - required angle in degrees is 90

OpenStudy (anonymous):

my fault it is pi/2.

OpenStudy (cwrw238):

as directrix said sine pi/2 = 1

OpenStudy (anonymous):

I know...*brain fart*